# 3D Optimization for AI Inference Scaling: Balancing Accuracy, Cost, and Latency

### v2

@@ Abstract
AI inference scaling is often tuned through 1D heuristics (a fixed reasoning passes) or 2D bivariate trade-offs (e.g., performance vs. compute), which fail to consider cost and latency constraints. We introduce a 3D optimization framework that jointly calibrates accuracy, cost, and latency within a unified decision space, enabling constraints-aware inference scaling. Using Monte Carlo simulations across three representative scenarios and nine simulated large language models, we evaluate four optimization methods to address the 3D multi-objective optimization (MOO) problem. Framing inference scaling in MOO shapes a feasible space that 1D and 2D optimizations fail to capture, enabling environment-adaptive selection of the inference scaling $k$. Results show that knee-point optimization achieves the best balance, while accuracy-maximization remains favorable when precision is prioritized. The framework establishes a theoretical foundation for deployment-aware inference scaling across diverse operational contexts. Code and simulation available at https://anonymous.4open.science/r/inf˙opt-FE40/README.md.
@@ Introduction
Beyond scaling AI training, which enhances performance by providing more data, scaling inference—which executes multiple reasoning paths before generating a final output—has been highlighted as an efficient method to improve performance (Snell et al. 2024) and reduce erroneous generations (Park et al. 2025). By sampling multiple responses, instead of relying on a single generation, inference scaling increases the likelihood of obtaining higher-quality outputs at the inference stage. This enables even smaller models to surpass the larger models (Wu et al. 2024; Wang et al. 2022). This approach is particularly efficient when parallel computing is available, where advances in AI infrastructure allow multiple inference instances to run concurrently (Bian, Yan, and Venkataraman 2025; NVIDIA 2025).
Figure 1: AI inference scaling optimization requires addressing the multidimensional tradeoff among cost, latency, and accuracy rather than relying on a bivariate tradeoff between performance-compute tradeoff that neglects constraints imposed by cost and latency in the deployment
Figure 2: Unlike conventional inference, which performs a single stochastic forward pass (k=1k{=}1) to generate one output, parallel inference scaling executes multiple inference batches concurrently ($k\geq 2$) and aggregates their outputs (e.g., best-of-$k$), thereby improving performance with minimal increase in latency. Parallelism is particularly effective when sufficient computational resources are available to support concurrent execution.
Prior research on AI inference scaling has primarily focused on 2D relationships between performance and inference scale, demonstrating accuracy gains with increased computational allocation (Wang et al. 2022; Jin, Wei, and Brooks 2025). This approach addresses the limitation of static inference scaling by allowing adaptive compute allocation. For example, in medical AI applications, users may willingly invest additional computational resources to obtain more reliable results. Executing multiple inference passes substantially increases the likelihood of generating higher-quality outputs compared to the single-shot inference (Snell et al. 2024). Such bivariate optimization effectively identifies the inference scale that maximizes accuracy under a limited compute budget. However, the 2D optimization fails to account for constraints on time and cost, factors that critically considered in the real-world deployment settings (NVIDIA 2025). Neglecting these dimensions renders existing scaling analyses incomplete for practical, deployment-aware optimization of AI inference.
@@
Inference scaling methods enhance model performance by generating multiple candidate outputs and selecting the best among them. Techniques such as best-of-$k$ and self-consistency demonstrate that additional sampling at inference time can enhance reasoning capacity without retraining (Sadhukhan et al. 2025; Wang et al. 2022). Parallelism makes this attainable by executing multiple inference passes simultaneously with minimal latency overhead, leveraging distributed compute resources such as GPUs (NVIDIA 2025). Recent studies also suggest that smaller models can match or even exceed the accuracy of larger ones when parallel inference is effectively utilized (Snell et al. 2024). While these techniques highlight the potential of scaling inference through parallelism, most studies focus on increasing performance via more computing, rather than configuring the optimality with both cost and latency.
@@
Another research line emphasized the need to incorporate cost and latency factors for real-world AI systems (NVIDIA 2025; Aubrey 2025). These studies highlight that inference performance is bounded by operational factors such as per-query cost ceilings and service-level latency requirements. Under such constraints, balancing performance, latency, and cost is crucial to achieve effective AI infrastructure operation  (Harris 2025). Most of the research remains grounded in bivariate performance-compute tradeoff curve (Brown et al. 2024; Snell et al. 2024; Wang et al. 2022; Jin, Wei, and Brooks 2025; Park et al. 2025). To the best of our knowledge, no prior research has formulated inference optimization that simultaneously accounts for cost, latency, and performance, despite increasing recognition of its necessity.
@@
Figure 3: Compute-optimal AI requires scaling efficiency in both training and inference. Training scaling laws (e.g., Chinchilla) guide how models are built, while inference scaling laws govern how they are used efficiently under real-world cost and latency constraints.
@@
The “best-of-k” is the most common AI inference strategy (Snell et al. 2024).
@@
Figure 4: Inference-scaling optimization results for GPT-5 across three simulated scenarios. (a–c) 3D feasible cubes in cost–time–accuracy space with constraint planes at $C_{\max}$, $T_{\max}$, and $A_{\min}$. Markers denote optimization criteria ($\bullet$ Accuracy-Optimal, $\blacktriangle$ Cube-Optimal, $\blacklozenge$ Utopia-Closest, $\blacksquare$ Knee-Point); colors indicate cube volume (larger is better), and the yellow star marks the utopia point.
@@
Figure 5: Accuracy-inference compute (K) trade-offs across three scenarios. These show Monte Carlo means with CI = 95%; the red dashed line indicates $A_{\min}$. Markers correspond to the optimal configurations identified in Fig. 4. Distinct operational priorities yield different optimal inference scales $k^{\star}$.
@@ Discussion
From 1D and 2D to 3D Optimization
Conventional approaches to inference scaling typically relied on 1D heuristics, such as executing a fixed number of reasoning passes (e.g., 16 passes), or on 2D tradeoffs between performance–scale. While these methods improved performance, they failed to capture the constraints of latency and cost that are essential to operation in AI infrastructure. The proposed 3D MOO overcomes the limitation of past approaches by jointly optimizing accuracy, cost, and latency. This perspective reveals new feasible regions that 1D and 2D optimization fail to capture, providing an opportunity for a deployment-aware inference optimization.
@@
The results demonstrate that additional compute allocation can compensate the performance for the smaller size. When parallelism is available, smaller models can achieve comparable or even superior accuracy to larger architectures with a lower cost and latency. For example, GPT-5 Nano and Nemotron 9B both matched the performance of their larger counterparts while operating under significantly lower latency and cost. These findings suggest that inference scaling offers a practical path toward the compute efficiency, especially in resource-constrained environments.
The Role of Knee-Point Optimization
Among MOOs, knee-point selection consistently achieved the most balanced efficiency in the 3D space. It identifies the point of maximum curvature on the Pareto frontier, showing the relative efficiency in tri-objectives and a pragmatic equilibrium between performance and efficiency. While it could be fail to be identified for the linearity of the frontiers that does not show the knee point, it can serve as the most practical solution except for the case that needs the maximal accuracy.
@@
This study relies on stochastic simulation rather than live LLM deployments, and thus abstracts away system-level factors such as GPU utilization, memory contention, and network queuing delays. Real-world AI infrastructures often exhibit non-Gaussian variability due to heterogeneous hardware, fluctuating energy conditions, and network overhead, all of which complicate inference-time optimization. Incorporating empirical traces from production environments—particularly latency and cost data from modern GPU clusters—would further validate and refine the proposed framework under realistic serving conditions.
Future work will extend this framework to infrastructure-aware optimization. This includes integrating hardware metrics from GPUs and Kubernetes-based orchestration to align inference scaling with real-time resource allocation. The optimization layer could operate as an adaptive router that dynamically selects both the model family and inference scale based on complexity and system load. Such integration would enable restriction and resource-aware optimization across heterogeneous model pools, improving inference efficiency in large-scale AI system operation.
@@ Conclusion
We reframe AI inference scaling as a multidimensional optimization problem that considers accuracy, cost, and latency in one decision space. By integrating Monte Carlo simulation with multiobjective optimization, the proposed framework bridges the gap between theoretical scaling laws and deployment-aware scaling. Across diverse model families and budget constraints, the result revealed that knee-optimality from the Pareto frontiers makes the scale that shows a relative efficiency among others. Beyond empirical findings, this framework establishes a foundation for advanced inference systems that respond to operational conditions in real time. Embedding MOO into AI routers and orchestration layers will enable efficient inference across heterogeneous AI models and diverse environments.
@@ Acknowledgments
Thanks to Nashua Springberry and Michael Schuler for constructive comments on the design and programming for the simulation.

### v3

@@ Abstract
AI inference scaling is often tuned through 1D heuristics (a fixed reasoning pass) or 2D bivariate trade-offs (e.g., accuracy vs. compute), which fail to consider cost and latency constraints. We introduce a 3D optimization framework that jointly calibrates accuracy, cost, and latency within a unified decision space, enabling constraints-aware inference scaling. Using Monte Carlo simulations across three representative scenarios and nine simulated large language models, we evaluate four optimization methods to address the 3D multi-objective optimization (MOO) problem. Framing inference scaling in MOO shapes a feasible space that 1D and 2D optimizations fail to capture, enabling environment-adaptive selection of the inference scaling $k$. Results show that knee-point optimization based on Pareto frontiers achieves the best balance, while accuracy-maximization remains favorable when accuracy is prioritized. Our results further show that smaller models, when combined with optimal inference scaling, can match or exceed the performance of larger models at a fraction of the cost. The framework establishes a theoretical foundation for deployment-aware inference scaling across diverse operational conditions. Code and simulation available at https://github.com/masonjung/inference-scaling-moo.
@@ Introduction
Beyond scaling AI training, which enhances performance by providing more data, scaling inference—which executes multiple reasoning paths before generating a final output—has been highlighted as an efficient method to improve performance (snell2024scaling) and reduce erroneous generations (park2025know). By sampling multiple responses, instead of relying on a single generation, inference scaling increases the likelihood of obtaining higher-quality outputs at the inference stage. This enables even smaller models to surpass the larger models (wu2024inference; wang2022self). This approach is particularly efficient when parallel computing is available, where advances in AI infrastructure allow multiple inference instances to run concurrently (bian2025scaling; nvidia2025balancing).
Figure 1: AI inference scaling requires optimizing the multidimensional tradeoff among cost, latency, and accuracy rather than relying on a bivariate performance-compute tradeoff that neglects the cost and latency constraints inherent in real-world deployment.
Figure 2: Unlike conventional inference, which performs a single stochastic forward pass (k=1k{=}1) to generate one output, parallel inference scaling executes multiple inference batches concurrently ($k\geq 2$) and aggregates their outputs (e.g., best-of-$k$) to improve accuracy with minimal increase in latency. Parallelism is particularly effective when sufficient computational resources are available to support concurrent execution.
Prior research on AI inference scaling has primarily focused on 2D relationships between performance and inference scale, demonstrating accuracy gains with increased computational allocation (wang2022self; jin2025energy). This approach addresses the limitation of static inference scaling by allowing adaptive compute allocation. For example, in medical AI applications, users may willingly invest additional computational resources to obtain more reliable results. Executing multiple inference passes substantially increases the likelihood of generating higher-quality outputs compared to the single-shot inference (snell2024scaling). Such bivariate optimization effectively identifies the inference scale that maximizes accuracy under a limited compute budget. However, the 2D optimization fails to account for constraints on time and cost, factors that critically considered in the real-world deployment settings (nvidia2025balancing). Neglecting these dimensions renders existing scaling analyses incomplete for practical, deployment-aware optimization of AI inference.
@@
Inference scaling methods enhance model performance by generating multiple candidate outputs and selecting the best among them. Techniques such as best-of-$k$ and self-consistency demonstrate that additional sampling at inference time can enhance reasoning capacity without retraining (sadhukhan2025kinetics; wang2022self). Parallelism makes this attainable by executing multiple inference passes simultaneously with minimal latency overhead, leveraging distributed compute resources such as GPUs (nvidia2025balancing). Recent studies also suggest that smaller models can match or even exceed the accuracy of larger ones when parallel inference is effectively utilized (snell2024scaling). While these techniques highlight the potential of scaling inference through parallelism, most studies focus on increasing performance via more computing, rather than configuring the optimality with both cost and latency.
@@
Another research line emphasized the need to incorporate cost and latency factors for real-world AI systems (nvidia2025balancing; aubrey2025maximizeai). These studies highlight that inference performance is bounded by operational factors such as per-query cost ceilings and service-level latency requirements. Under such constraints, balancing performance, latency, and cost is crucial to achieve effective AI infrastructure operation  (harris2025thinksmart). Most of the research remains grounded in bivariate performance-compute tradeoff curve (brown2024large; snell2024scaling; wang2022self; jin2025energy; park2025know). To the best of our knowledge, no prior research has formulated inference optimization that simultaneously accounts for cost, latency, and performance, despite increasing recognition of its necessity.
@@
Figure 3: Compute-optimal AI system requires scaling efficiency in both training and inference. Training scaling laws (e.g., Chinchilla law) guide how models are built, while inference scaling laws govern how they are used efficiently under real-world cost and latency constraints.
@@
The “best-of-k” is the most common AI inference strategy (snell2024scaling).
@@
Figure 4: Inference-scaling optimization results for GPT-5 across three simulated scenarios. (a–c) 3D feasible cubes in cost–time–accuracy space with constraint planes of maximum cost, maximum latency, and minimal ACC ($C_{\max}$, $T_{\max}$, and $A_{\min}$). Markers denote each optimal point ($\bullet$ Maximum-ACC, $\blacktriangle$ Cube-Optimal, $\blacklozenge$ Utopia-Closest, $\blacksquare$ Knee-Point); purple colors indicate cube volume (larger is better), and the yellow star marks the utopia point.
@@
Figure 5: Accuracy-inference scale (k) trade-offs across three scenarios for GPT-5. These show MC means with CI = 95%; the red dashed line indicates minimal ACC ($A_{\min}$). Markers correspond to the optimal configurations identified in Fig. 4. Distinct operational priorities yield different optimal inference scales $k^{\star}$.
@@ Discussion
Toward 3D Optimization from 1D and 2D
Conventional approaches to inference scaling typically relied on 1D heuristics, such as executing a fixed number of reasoning passes (e.g., 16 passes), or on 2D tradeoffs between performance and compute scale. While these methods improved accuracy, they failed to capture the constraints of latency and cost that are essential for AI inference. Because real-world inference should operate within given cost and latency budgets, ignoring these constraints leads to impractical inference. The proposed 3D MOO overcomes the limitation of past approaches by jointly optimizing accuracy, cost, and latency. This method reveals new feasible regions that 1D and 2D optimization fail to capture. The 3D objective space provides an opportunity for a deployment-aware inference optimization.
@@
The results demonstrate that additional compute allocation can compensate the performance for the smaller size. When parallelism is available, smaller models can achieve comparable or even superior accuracy to larger architectures with a lower cost and latency. For example, GPT-5 Nano with 56 inference paths showed ACC of 0.99 (\$0.05, 30.1s), which is greater than GPT-5 with 24 inference paths that achieved ACC of 0.98 (\$0.52, 64.5s) under the knee-point selection criterion in Scenario 3. This result suggests that inference scaling offers an efficient path for small models, especially when parallel inference is available.
Knee-Point Optimization for the Best Balance
Among MOOs, knee-point selection consistently achieved the most balanced efficiency in the feasible space. The Knee MOO identifies the point of maximum curvature on the Pareto frontier, showing the relative efficiency in tri-objectives. While it could be fail to be identified for the linearity of the frontiers that does not show the knee point, it can serve as the most practical solution except for the case that needs the maximal accuracy. Regarding the inference for the minimal cost and the best latency, the solution is always inference with a single path.
@@
This study relies on stochastic simulation rather than LLM deployments. Thus, it abstracts away system-level factors such as GPU utilization, memory contention, and network queuing delays. Real-world AI infrastructures often exhibit non-Gaussian variability due to heterogeneous hardware, fluctuating energy conditions, and network overhead, all of which complicate inference MOO for further dimensions and higher dimensions. Incorporating empirical traces from AI infrastructure environments would further validate the proposed framework.
Also, the proposed AI inference MOO framework is critical not only for single-agent models but also for multi-agent systems, where multiple agents and tools compound the cost and latency overhead, further amplifying the need for efficient inference scaling. Moreover, as dimensions such as energy consumption, safety constraints, and thermal limits become increasingly relevant, the framework can be extended beyond 3D to support higher-dimensional MOO. Such a generalized formulation would enable jointly optimizing energy, safety, and heat dissipation on top of the proposed 3D space will support deployment-aware reasoning for real-world AI systems.
Future work could extend our framework to infrastructure-aware optimization. This includes integrating hardware metrics from GPUs and Kubernetes-based orchestration to align inference scaling with real-time resource allocation. The optimization layer could operate as an adaptive router that dynamically selects the model and inference scale based on complexity and system load. Such integration would enable real-world MOO across heterogeneous pools, improving inference efficiency on a large scale.
@@ Conclusion
We reframe AI inference scaling as a multidimensional optimization problem that considers accuracy, cost, and latency in one optimization to overcome the limitations of 1D and 2D optimization. By integrating MC simulations with MOO, the proposed framework enables constraint-aware inference scaling. Across diverse model families and scenarios, the result revealed that knee-optimality from the Pareto frontiers makes the scale that shows a relative efficiency among others. Also, we found that smaller models could show a better performance than larger models with a lower cost and latency. Beyond these findings, this framework establishes a foundation for advanced inference systems that respond to operational conditions. Embedding MOO into orchestration layers will enable efficient inference across heterogeneous AI models and diverse environments. As foundation models and agentic systems continue to expand, principled inference scaling will become essential to deploy AI efficiently.
@@ Acknowledgments
This project is a part of Cloudera’s Accelerators for ML Projects (AMPs). The AMP is accessible via https://github.com/cloudera/CAI˙AMP˙Inference˙Scaling˙Optimization. Thanks to Nashua Springberry and Michael Schuler for constructive comments on the design and programming for the simulation.